OpenStudy (anonymous):

Cork has a density of 0.240 g/cm´. What volume would 10.0 grams of cork have?

OpenStudy (anonymous):

the density will not change. that is, the ratio of mass to volume will be a constant. i.e., mass / volume = 0.240 grams / cm^3 = 10 grams / x cm^3

OpenStudy (shinalcantara):

\[\rho = \frac{ m }{ V }\] where rho is the density m is mass V is volume getting the volume you'll have: \[V = \frac{ m }{ \rho }\] substituting the values \[V = \frac{ 10g }{ 0.24g/cm^3 }\] solve for the volume